A production lot of size 100 is known to be a% defective. A random sample of 12

items is selected without replacement. Determine the probability that there will be

no defective in the sample.


1
Expert's answer
2021-08-11T07:05:27-0400

Let "X=" the number of defective items in the sample: "x\sim Bin(n, p)."

Given "n=12, p=0.01a, q=1-p=1-0.01a"


"P(X=0)=\dbinom{12}{0}(0.01a)^0(1-0.01a)^{12-0}"

"=(1-0.01a)^{12}"